OpenStudy (anonymous):

y = log x If y = 10, then what is x?

OpenStudy (nincompoop):

what do you know about logs?

OpenStudy (anonymous):

The logarithm of a number is the exponent by which another fixed value, the base, must be raised to produce that number. For example, the logarithm of 1000 to base 10 is 3, because 1000 is 10 to the power 3: 1000 = 10 × 10 × 10 = 103. More generally, if x = by, then y is the logarithm of x to base b, and is written y = logb(x), so log10(1000) = 3.

OpenStudy (nincompoop):

nice copy and paste of a definition, dude what do you understand about logs?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

lol

OpenStudy (nincompoop):

well we can try to understand your definition as a start

OpenStudy (anonymous):

10=10^x

OpenStudy (anonymous):

but is this the right way

OpenStudy (noelgreco):

When you see a logarithm like y=log x ask yourself the question, "to what power must I raise 10 (the base) to get x?" If y is 10 then x is 10^10.

OpenStudy (nincompoop):

\[2\times2\times2=(2)^3=8\] In this case, your logarithm is 3 and we can write it like this \[\log_{2} (8)=3\]

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

i have to solve for x

OpenStudy (nincompoop):

the number we are multiplying is called base, so we can say that: the logarithm of base 2 is 3

OpenStudy (nincompoop):

I am providing you the fundamental concept to logarithms

OpenStudy (anonymous):

nvm i will google it

OpenStudy (nincompoop):

you will get the same concept as what I am teaching you now

OpenStudy (nincompoop):

http://www.purplemath.com/modules/logs.htm

OpenStudy (anonymous):

ok so i get it now so its y = log x 10 = log x 10 = 10^x x = 1

OpenStudy (agent0smith):

@RAWFECTION y = log x 10 = log x 10 = 10^x x = 1 this is not done correctly. x is not equal to 1... log1 equals zero, not 10.

OpenStudy (agent0smith):

Correct solution is: \[ \log_{10} x = 10\] so \[\large 10^{10} = x\] ^that is x. This comes from this rule: \[\large \log_{a} x = b\] then \[\large a^b = x\]